# PADiff: Predictive and Adaptive Diffusion Policies for Ad Hoc Teamwork

### v1

@@ Introduction
Ad hoc teamwork (AHT) presents a core challenge in multi-agent systems (stone2010ad), which requires agents to collaborate with unknown teammates without predefined coordination (yuan2023learning). Consider a robotic soccer match, where an autonomous agent plays with unfamiliar teammates. The agent must make real-time decisions without any knowledge of the new teammates’ playing styles or strategies. These challenges are not limited to simulated environments. In real-world applications like disaster response (yucesoy2025role), agents must coordinate with human partners operating under unknown procedures. In autonomous driving (teng2023motion), vehicles often need to interact with other road users without established communication protocols. Such situations require agents to infer teammate behaviors, adapt their strategies, and cooperate effectively on the fly (ravula2019ad). As multi-agent systems are deployed in real world (yuan2023survey), it is urgent to address the challenges of AHT and improve the autonomy of multi-agent systems.
@@
AHT fundamentally requires an ego agent to collaborate with previously unseen teammates, making it crucial to prepare for multiple possible cooperative patterns rather than relying on a single optimal response. Consider a robotic soccer scenario involving AHT shown in Figure 1): When an ego agent advances with the ball, she has to choose between two modes of cooperation: either assist a teammate or directly attempt a shot. This example indicates that the ego agent must learn multiple cooperation patterns simultaneously, highlighting the multimodal nature of the optimal policies. However, most existing solutions to AHT are based on reinforcement learning (RL) (chen2020aateam; gu2021online; papoudakis2021agent), which often fail to learn multimodal policies. RL typically optimizes for a fixed expected return, causing the policy to collapse to a single dominant behavior. From the generalization perspective, even if RL learns an optimal policy, it often still fails to generalize when the teammates change their cooperation patterns. Although previous AHT methods (e.g., LIAM(papoudakis2021agent) and GPL-SPI(rahman2021towards)) leverage maximum entropy RL (e.g., SAC) to enhance stochastic exploration and mitigate policy collapse into a singular cooperative mode, their reliance on undirected behavioral dispersion inherently fails to structurally model multimodal distributions, thus poorly capturing diverse cooperation strategies. In contrast, diffusion models (croitoru2023diffusion) are naturally suitable to model multimodal distributions(wang2022diffusion), making them stronger potential solutions to AHT, where learning multiple cooperation patterns simultaneously is essential for robust teamwork.
@@
To address these two limitations, we present PADiff, a diffusion-based framework for learning Predictive and Adaptive policies in AHT. The core innovations of PADiff are as follows. (1) To address the lack of predictive capability in traditional diffusion models, we design a Predictive Guidance Block (PGB) integrated into the denoising process. This module specifically leverages intermediate representations to predict teammates’ cooperative targets and align action generation with long-term team objectives. (2) To address the limited adaptability of traditional diffusion architectures, we design an Adaptive Feature Modulation Net (AFM-Net), which integrates two FiLM-like feature-wise modulation layers to scale and shift intermediate features based on the current state (perez2018film). To further improve the diffusion policies’ stability and adaptability against uncertain teammates, AFM-Net integrates several designs such as Layer Normalization, Residual connections and Dropout regularization into the training process. Through this way, AFM-Net dynamically adjusts its internal representations to accommodate the frequently changing teammates’ goals and behaviors.
@@
Ad Hoc Teamwork. The core challenge of AHT is to develop a policy that generalizes across diverse teammate behaviors and enables rapid adaptation during execution. Early approaches typically rely on predefined sets of teammate types, using type inference to switch among corresponding policies (e.g., PLASTIC (barrett2015cooperating)). While effective in constrained environments, these methods (chen2020aateam; barrett2015cooperating; durugkar2020balancing; mirsky2020penny) struggle when teammates exhibit complex or unmodeled behaviors. More recent works (papoudakis2021agent; gu2021online) improve adaptability by learning representations of teammates, allowing the agent to adjust its policy based on online interactions. These methods relax the assumption of fixed types and broaden the space of potential collaborations. However, most existing approaches (barrett2017making; chen2020aateam; rahman2021towards; zhang2025ad) are still built on RL frameworks, where the optimization objective focuses on maximizing expected return. This leads the learned policy to converge toward a single high-reward action, neglecting the diversity of possible cooperative modes under different teammate dynamics. While maximum entropy regularization ((haarnoja2018soft)) mitigates this collapse by promoting exploratory behavior, it fundamentally fails to construct genuinely multimodal policies. Such limited policy expressiveness directly constrains generalization in AHT. In contrast, diffusion models (zhu2023diffusion) have the ability to represent any normalizable distribution, with the potential to improve the AHT policy representation effectively.
@@
Generative Models for Decision Making. Generative Models for Decision Making. Generative models have seen increasing use in decision-making tasks (gozalo2023chatgpt; jo2023promise; brynjolfsson2025generative; goodfellow2020generative). While GAN-based methods (goodfellow2020generative) like GAIL (ho2016generative) offer adversarial imitation, they often struggle with training instability and limited generalization. VAEs (kingma2013auto; mandlekar2020learning; mees2022matters; lynch2020learning) utilize latent representations for policy generation but face difficulties capturing fine-grained dynamics. More recently, Transformer-based models (chen2021decision; zheng2022online; xie2023future), notably Decision Transformer (chen2021decision), have reframed policy learning as return-conditioned sequence modeling, achieving strong offline RL performance. However, most generative policy models remain unimodal due to architectural biases like Gaussian priors and autoregressive decoding. In contrast, diffusion-based policies (wang2022diffusion) can learn multimodal action distributions. Leveraging this, we embed teammate-predictive guidance into a diffusion-based policy framework specifically for AHT.
@@
Diffusion Models. Diffusion models (sohl2015deep), first for general high-dimensional data, rapidly found success in image synthesis. Their strength here lies in U-Net backbones’ multi-scale feature aggregation and spatial inductive bias (ho2020denoising; nichol2021improved; vahdat2021score). DiT (peebles2023scalable) replaces the U-Net backbone with Vision Transformers, showing that attention mechanism improves sample quality by capturing long-range dependencies. Beyond vision, diffusion models have recently been adopted for decision-making in RL field. Trajectory-level planners (janner2022planning; liang2023adaptdiffuser; liang2024skilldiffuser; zhang2024motiondiffuse) generate full state–action sequences for long-horizon reasoning, typically with U-Nets (janner2022planning) or Transformer variants like DiT (zhang2024motiondiffuse). Action-level policies, by contrast, denoise one action at a time with lightweight MLPs (wang2022diffusion; kang2023efficient; hansen2023idql), enabling real-time control. We follow the action-level paradigm but tackle the more demanding AHT setting. To this end, we introduce (i) a Conditional Feature Modulation Network that sustains robustness under dynamically changing teammate policies and (ii) a Predictive Guidance Block that injects teammate-aware goals into the denoising process.
@@
Diffusion Probabilistic Models. Diffusion models (DMs) (sohl2015deep; song2019generative) are a powerful type of generative models that learn to recover data samples from noise. This typically involves two phases: a forward noising process that gradually perturbs clean data into noise, and a reverse process that learns to reconstruct data from noise. Given a data sample $x_{0}\sim p_{\text{data}}(x)$, the forward process defines a Markov chain: $p(x_{k}\mid x_{k-1})=\mathcal{N}(x_{k}\mid\sqrt{\alpha_{k}}x_{k-1},(1-\alpha_{k})I),$ where $\alpha_{k}$ controls the noise level at step $k$. The reverse process is parameterized by: $q_{\theta}(x_{k-1}\mid x_{k})=\mathcal{N}(x_{k-1}\mid\mu_{\theta}(x_{k},k),(1-\alpha_{k})I).$ Sampling starts from $x_{K}\sim\mathcal{N}(0,I)$ and iteratively denoises toward $x_{0}$. The model is trained by minimizing the loss:
@@ Methods
For discrete action spaces, following D3PM (austin2021structured), the reverse diffusion is parameterized as:
@@
where $\overline{Q}_{k}=Q_{1}\cdot Q_{2}\dots Q_{k}$ and the schedule type of transition matrices $Q$ follows the Uniform approach, which is a way of controlling the forward diffusion process and the learnable reverse diffusion process. $\text{Cat}(\bm{x};\bm{p})$ is a categorical distribution over the one-hot row vector $\bm{x}$ with probabilities given by the row vector $\bm{p}$ (austin2021structured).
@@
State Encoder: To better identify the cooperation situation of the team and handle potential changes in uncertain teammates, we model the team’s cooperation state using a history window $s_{t-m:t}$ which provides richer context about interactions. This sequence is compressed into an informative latent space modeled as a multivariate Gaussian distribution (gu2021online). A sample from this latent distribution acts as the conditional input during the denoising stage.
@@
Adaptive Feature Modulation Net (AFM-Net): Traditional denoising networks, such as UNet and MLP, lack necessary teammate-awareness, while Transformer-based networks like DiT(peebles2023scalable) rely on attention mechanisms that introduce significant computational overhead, undesirable in fast-paced AHT settings. To overcome these limitations, we propose AFM-Net, an effective and robust denoising network that integrates the FiLM mechanism, residual connections, layer normalization and dropout regularization, making it both expressive and efficient without the need for attention. AFM-Net has three primary features: (1) Conditional feature modulation through FiLM-style(perez2018film) method, enhancing teammate adaptability, (2) Residual connections for stable training and robust representations and (3) Dropout regularization to improve generalization to unseen teammates.
@@
Unlike conventional inference-time guidance methods, which often lack flexibility in adapting to real-time teammate dynamics, PGB leverages the intermediate representation generated during the denoising process to predict teammates’ intentions. By propagating gradients through the denoising process, it optimizes this representation, enabling it to anticipate teammates’ behaviors. As a result, the denoising process becomes adaptive to changes in teammates’ actions. The design of the PGB contains two core predictive tasks: Collaborative Return (CoReturn) represents the expected cumulative future team reward, conditioned on the current state and intermediate feature from the denoising network. The prediction of CoReturn serves to guide the generation of actions towards better alignment with overall team objectives. And Collaborative Goal (CoGoal) stands for the predicted future state, which is served as a sub-goal to capture the intention of teammates. While model-base method like DreamerV3(hafner2023mastering) optimize policies through environment prediction, PGB similarly enhances decision-making via auxiliary prediction tasks, but fundamentally differs by tailoring predictions to real-time collaborative intent. Each predictive task contributes distinctively to enhanced robust cooperative decision-making.
@@
Collections of teammates policies. In AHT scenarios, we must ensure that the ego agent can adapt to previously unseen teammates. To build a diverse pool of teammate policies, we use the Soft-Value Diversity (SVD) method from the CSP(ding2023coordination) framework to collect teammate behaviors, instantiating four independent multi‑agent populations in each environments. The training process alternates between the inner loop and the outer loop. In the inner loop, each team is trained in isolation, updating its policy to maximize its own performance. In the outer loop, we jointly update all teams to maximize the SVD objective, encouraging them to develop distinct value estimates over observation–action pairs, thus diversifying their behaviors. Specifically, three training populations are dedicated to interacting with the ego agent for collecting the training data, while a fourth population is reserved exclusively for assessing the ego agent’s performance. Each training population comprises three unique policy checkpoints, whereas the testing population includes twelve checkpoints. These checkpoints correspond to distinct joint policies. By isolating the training and testing in this way, our protocol ensures a rigorous evaluation of the model’s ability to adapt to previously unseen teammate strategies. We divided the 12 test policies into two groups, one containing 4 test strategies and the other containing 8. The cross-play matrices experiments reveal that cooperation scores for different teammate combinations are lower than for identical teammate combinations. This is due to strategy differences causing coordination issues, which reflects the diversity of test teammates strategies. Detailed information about cross-play experiments can be found in Appendix A.2.
Baselines. In our experiments, we compare two classic AHT methods: LIAM (papoudakis2021agent) and ODITS (gu2021online). To underscore the benefits of our method and its superior adaptability to AHT tasks, we further benchmark it against two diffusion‑based models: Diffusion‑BC and Diffusion‑QL (wang2022diffusion). We also compare with the offline reinforcement learning algorithm CQL (kumar2020conservative) and two AHT–focused multi‑agent methods: MADIFF (zhu2024madiff), which aggregates all agents’ observations into a global state sequence and uses an inverse dynamics model to predict the ego agent’s next action; and MADT (meng2021offline), which only trains the ego agent’s actor network within the AHT settings. We also take the latest DT-based offline AHT method TAGET (zhang2025ad) as baseline. These comparisons demonstrate that our method addresses the challenges in AHT scenarios more effectively and achieve significant performance.

### v2

@@ Introduction
Ad hoc teamwork (AHT) presents a core challenge in multi-agent systems (Stone et al. 2010), which requires agents to collaborate with unknown teammates without predefined coordination (Yuan et al. 2023a). Consider a robotic soccer match, where an autonomous agent plays with unfamiliar teammates. The agent must make real-time decisions without any knowledge of the new teammates’ playing styles or strategies. These challenges are not limited to simulated environments. In real-world applications like disaster response (Yucesoy et al. 2025), agents must coordinate with human partners operating under unknown procedures. In autonomous driving (Teng et al. 2023), vehicles often need to interact with other road users without established communication protocols. Such situations require agents to infer teammate behaviors, adapt their strategies, and cooperate effectively on the fly (Ravula 2019). As multi-agent systems are deployed in real world (Yuan et al. 2023b), it is urgent to address the challenges of AHT and improve the autonomy of multi-agent systems.
@@
AHT fundamentally requires an ego agent to collaborate with previously unseen teammates, making it crucial to prepare for multiple possible cooperative patterns rather than relying on a single optimal response. Consider a robotic soccer scenario involving AHT shown in Figure 1): When an ego agent advances with the ball, she has to choose between two modes of cooperation: either assist a teammate or directly attempt a shot. This example indicates that the ego agent must learn multiple cooperation patterns simultaneously, highlighting the multimodal nature of the optimal policies. However, most existing solutions to AHT are based on reinforcement learning (RL) (Chen et al. 2020; Gu et al. 2021; Papoudakis et al. 2021), which often fail to learn multimodal policies. RL typically optimizes for a fixed expected return, causing the policy to collapse to a single dominant behavior. From the generalization perspective, even if RL learns an optimal policy, it often still fails to generalize when the teammates change their cooperation patterns. Although previous AHT methods (e.g., LIAM(Papoudakis et al. 2021) and GPL-SPI(Rahman et al. 2021)) leverage maximum entropy RL (e.g., SAC) to enhance stochastic exploration and mitigate policy collapse into a singular cooperative mode, their reliance on undirected behavioral dispersion inherently fails to structurally model multimodal distributions, thus poorly capturing diverse cooperation strategies. In contrast, diffusion models (Croitoru et al. 2023) are naturally suitable to model multimodal distributions(Wang et al. 2022), making them stronger potential solutions to AHT, where learning multiple cooperation patterns simultaneously is essential for robust teamwork.
@@
To address these two limitations, we present PADiff, a diffusion-based framework for learning Predictive and Adaptive policies in AHT. The core innovations of PADiff are as follows. (1) To address the lack of predictive capability in traditional diffusion models, we design a Predictive Guidance Block (PGB) integrated into the denoising process. This module specifically leverages intermediate representations to predict teammates’ cooperative targets and align action generation with long-term team objectives. (2) To address the limited adaptability of traditional diffusion architectures, we design an Adaptive Feature Modulation Net (AFM-Net), which integrates two FiLM-like feature-wise modulation layers to scale and shift intermediate features based on the current state (Perez et al. 2018). To further improve the diffusion policies’ stability and adaptability against uncertain teammates, AFM-Net integrates several designs such as Layer Normalization, Residual connections and Dropout regularization into the training process. Through this way, AFM-Net dynamically adjusts its internal representations to accommodate the frequently changing teammates’ goals and behaviors.
@@
Ad Hoc Teamwork. The core challenge of AHT is to develop a policy that generalizes across diverse teammate behaviors and enables rapid adaptation during execution. Early approaches typically rely on predefined sets of teammate types, using type inference to switch among corresponding policies (e.g., PLASTIC (Barrett and Stone 2015)). While effective in constrained environments, these methods (Chen et al. 2020; Barrett and Stone 2015; Durugkar et al. 2020; Mirsky et al. 2020) struggle when teammates exhibit complex or unmodeled behaviors. More recent works (Papoudakis et al. 2021; Gu et al. 2021) improve adaptability by learning representations of teammates, allowing the agent to adjust its policy based on online interactions. These methods relax the assumption of fixed types and broaden the space of potential collaborations. However, most existing approaches (Barrett et al. 2017; Chen et al. 2020; Rahman et al. 2021; Zhang et al. 2025) are still built on RL frameworks, where the optimization objective focuses on maximizing expected return. This leads the learned policy to converge toward a single high-reward action, neglecting the diversity of possible cooperative modes under different teammate dynamics. While maximum entropy regularization ((Haarnoja et al. 2018)) mitigates this collapse by promoting exploratory behavior, it fundamentally fails to construct genuinely multimodal policies. Such limited policy expressiveness directly constrains generalization in AHT. In contrast, diffusion models (Zhu et al. 2023) have the ability to represent any normalizable distribution, with the potential to improve the AHT policy representation effectively.
@@
Generative Models for Decision Making. Generative Models for Decision Making. Generative models have seen increasing use in decision-making tasks (Gozalo-Brizuela and Garrido-Merchan 2023; Jo 2023; Brynjolfsson et al. 2025; Goodfellow et al. 2020). While GAN-based methods (Goodfellow et al. 2020) like GAIL (Ho and Ermon 2016) offer adversarial imitation, they often struggle with training instability and limited generalization. VAEs (Kingma et al. 2013; Mandlekar et al. 2020; Mees et al. 2022; Lynch et al. 2020) utilize latent representations for policy generation but face difficulties capturing fine-grained dynamics. More recently, Transformer-based models (Chen et al. 2021; Zheng et al. 2022; Xie et al. 2023), notably Decision Transformer (Chen et al. 2021), have reframed policy learning as return-conditioned sequence modeling, achieving strong offline RL performance. However, most generative policy models remain unimodal due to architectural biases like Gaussian priors and autoregressive decoding. In contrast, diffusion-based policies (Wang et al. 2022) can learn multimodal action distributions. Leveraging this, we embed teammate-predictive guidance into a diffusion-based policy framework specifically for AHT.
@@
Diffusion Models. Diffusion models (Sohl-Dickstein et al. 2015), first for general high-dimensional data, rapidly found success in image synthesis. Their strength here lies in U-Net backbones’ multi-scale feature aggregation and spatial inductive bias (Ho et al. 2020; Nichol and Dhariwal 2021; Vahdat et al. 2021). DiT (Peebles and Xie 2023) replaces the U-Net backbone with Vision Transformers, showing that attention mechanism improves sample quality by capturing long-range dependencies. Beyond vision, diffusion models have recently been adopted for decision-making in RL field. Trajectory-level planners (Janner et al. 2022; Liang et al. 2023, 2024; Zhang et al. 2024) generate full state–action sequences for long-horizon reasoning, typically with U-Nets (Janner et al. 2022) or Transformer variants like DiT (Zhang et al. 2024). Action-level policies, by contrast, denoise one action at a time with lightweight MLPs (Wang et al. 2022; Kang et al. 2023; Hansen-Estruch et al. 2023), enabling real-time control. We follow the action-level paradigm but tackle the more demanding AHT setting. To this end, we introduce (i) a Conditional Feature Modulation Network that sustains robustness under dynamically changing teammate policies and (ii) a Predictive Guidance Block that injects teammate-aware goals into the denoising process.
@@
Diffusion Probabilistic Models. Diffusion models (DMs) (Sohl-Dickstein et al. 2015; Song and Ermon 2019) are a powerful type of generative models that learn to recover data samples from noise. This typically involves two phases: a forward noising process that gradually perturbs clean data into noise, and a reverse process that learns to reconstruct data from noise. Given a data sample $x_{0}\sim p_{\text{data}}(x)$, the forward process defines a Markov chain: $p(x_{k}\mid x_{k-1})=\mathcal{N}(x_{k}\mid\sqrt{\alpha_{k}}x_{k-1},(1-\alpha_{k})I),$ where $\alpha_{k}$ controls the noise level at step $k$. The reverse process is parameterized by: $q_{\theta}(x_{k-1}\mid x_{k})=\mathcal{N}(x_{k-1}\mid\mu_{\theta}(x_{k},k),(1-\alpha_{k})I).$ Sampling starts from $x_{K}\sim\mathcal{N}(0,I)$ and iteratively denoises toward $x_{0}$. The model is trained by minimizing the loss:
@@ Methods
For discrete action spaces, following D3PM (Austin et al. 2021), the reverse diffusion is parameterized as:
@@
where $\overline{Q}_{k}=Q_{1}\cdot Q_{2}\dots Q_{k}$ and the schedule type of transition matrices $Q$ follows the Uniform approach, which is a way of controlling the forward diffusion process and the learnable reverse diffusion process. $\text{Cat}(\bm{x};\bm{p})$ is a categorical distribution over the one-hot row vector $\bm{x}$ with probabilities given by the row vector $\bm{p}$ (Austin et al. 2021).
@@
State Encoder: To better identify the cooperation situation of the team and handle potential changes in uncertain teammates, we model the team’s cooperation state using a history window $s_{t-m:t}$ which provides richer context about interactions. This sequence is compressed into an informative latent space modeled as a multivariate Gaussian distribution (Gu et al. 2021). A sample from this latent distribution acts as the conditional input during the denoising stage.
@@
Adaptive Feature Modulation Net (AFM-Net): Traditional denoising networks, such as UNet and MLP, lack necessary teammate-awareness, while Transformer-based networks like DiT(Peebles and Xie 2023) rely on attention mechanisms that introduce significant computational overhead, undesirable in fast-paced AHT settings. To overcome these limitations, we propose AFM-Net, an effective and robust denoising network that integrates the FiLM mechanism, residual connections, layer normalization and dropout regularization, making it both expressive and efficient without the need for attention. AFM-Net has three primary features: (1) Conditional feature modulation through FiLM-style(Perez et al. 2018) method, enhancing teammate adaptability, (2) Residual connections for stable training and robust representations and (3) Dropout regularization to improve generalization to unseen teammates.
@@
Unlike conventional inference-time guidance methods, which often lack flexibility in adapting to real-time teammate dynamics, PGB leverages the intermediate representation generated during the denoising process to predict teammates’ intentions. By propagating gradients through the denoising process, it optimizes this representation, enabling it to anticipate teammates’ behaviors. As a result, the denoising process becomes adaptive to changes in teammates’ actions. The design of the PGB contains two core predictive tasks: Collaborative Return (CoReturn) represents the expected cumulative future team reward, conditioned on the current state and intermediate feature from the denoising network. The prediction of CoReturn serves to guide the generation of actions towards better alignment with overall team objectives. And Collaborative Goal (CoGoal) stands for the predicted future state, which is served as a sub-goal to capture the intention of teammates. While model-base method like DreamerV3(Hafner et al. 2023) optimize policies through environment prediction, PGB similarly enhances decision-making via auxiliary prediction tasks, but fundamentally differs by tailoring predictions to real-time collaborative intent. Each predictive task contributes distinctively to enhanced robust cooperative decision-making.
@@
Collections of teammates policies. In AHT scenarios, we must ensure that the ego agent can adapt to previously unseen teammates. To build a diverse pool of teammate policies, we use the Soft-Value Diversity (SVD) method from the CSP(Ding et al. 2023) framework to collect teammate behaviors, instantiating four independent multi‑agent populations in each environments. The training process alternates between the inner loop and the outer loop. In the inner loop, each team is trained in isolation, updating its policy to maximize its own performance. In the outer loop, we jointly update all teams to maximize the SVD objective, encouraging them to develop distinct value estimates over observation–action pairs, thus diversifying their behaviors. Specifically, three training populations are dedicated to interacting with the ego agent for collecting the training data, while a fourth population is reserved exclusively for assessing the ego agent’s performance. Each training population comprises three unique policy checkpoints, whereas the testing population includes twelve checkpoints. These checkpoints correspond to distinct joint policies. By isolating the training and testing in this way, our protocol ensures a rigorous evaluation of the model’s ability to adapt to previously unseen teammate strategies. We divided the 12 test policies into two groups, one containing 4 test strategies and the other containing 8. The cross-play matrices experiments reveal that cooperation scores for different teammate combinations are lower than for identical teammate combinations. This is due to strategy differences causing coordination issues, which reflects the diversity of test teammates strategies. Detailed information about cross-play experiments can be found in Appendix A.2.
Baselines. In our experiments, we compare two classic AHT methods: LIAM (Papoudakis et al. 2021) and ODITS (Gu et al. 2021). To underscore the benefits of our method and its superior adaptability to AHT tasks, we further benchmark it against two diffusion‑based models: Diffusion‑BC and Diffusion‑QL (Wang et al. 2022). We also compare with the offline reinforcement learning algorithm CQL (Kumar et al. 2020) and two AHT–focused multi‑agent methods: MADIFF (Zhu et al. 2024), which aggregates all agents’ observations into a global state sequence and uses an inverse dynamics model to predict the ego agent’s next action; and MADT (Meng et al. 2021), which only trains the ego agent’s actor network within the AHT settings. We also take the latest DT-based offline AHT method TAGET (Zhang et al. 2025) as baseline. These comparisons demonstrate that our method addresses the challenges in AHT scenarios more effectively and achieve significant performance.